# Cross-Subject Data Splitting for Brain-to-Text Decoding

### v3

@@ 2 Related Work
3 Methodology
In this section, we will first introduce the formal definition of brain-to-text decoding and the general description of dataset format. Then we systematically analyze current cross-subject data splitting methods and point out that all existing methods suffer from two kinds of data leakage issues: brain signal leakage and text stimuli leakage. Finally, a cross-subject splitting criterion is proposed which avoids the above-mentioned data leakage problems.
3.1 Task Definition
Given the brain signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, consistent text segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$, where $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$, and $|T_{j}|=|F_{ij}|=L$. To EEG, since signals corresponding to complete text stimuli are available and they are continuous, we bond text $T_{j}$ (i.e. text stimuli) and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most scenarios, each text stimulus $T_{j}$ belongs to one certain task $M_{k}$. So the signal-text pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
Figure 3: The process of our proposed cross-subject data splitting method. (Color printing is preferred.)
One of the purposes of cross-subject splitting is to endow models the ability to decode unseen subject’s brain signal. As a result, if brain signal $F_{ij}$ appears in test set $S_{test}$, any signal $F_{i*}$ belonging to subject $i$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. The above splitting rules for training set can be formulated by Cartesian product:
$S_{train}=F_{train}\times T_{train},$ (1)
where
$F_{train}=\{F_{ij}|i\in I\},$ (2)
$I=\{i|F_{ij}\notin S_{val},S_{test},\forall j\},$ (3)
and
$T_{train}=\{T_{kj}|T_{kj}\notin S_{val},S_{test}\}.$ (4)
Such rules are also applicable to validation and test set. We omit their displays here for simplicity.
3.2 Current Splitting Methods
Current cross-subject data splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, the five dataset splitting methods are characterised as (1) split by subjects $S_{i}$, (2) split by tasks $M_{k}$, (3) split by randomly picking signal frames $F_{ij}$, (4) split by randomly picking signal frames under certain task $F_{ij}$-$M_{k}$, (5) split by randomly picking consecutive signal frames under certain task $F_{ij}$-$M_{k}$, corresponding to (a), (b), (c), (d), (e) in Figure 2. Figure 2 vividly displays the differences between current dataset splitting methods. In this example, we choose 4 subjects ($S_{1}$ to $S_{4}$) with 3 tasks ($M_{1}$ to $M_{3}$) each containing 4 ($T_{11}$ to $T_{14}$), 3 ($T_{21}$ to $T_{23}$), 4 ($T_{31}$ to $T_{34}$) text stimuli respectively. $F_{ij}$ indicates the brain signal of $i$-th subject stimulated by $j$-th text under task $M_{k}$, e.g. $F_{21}$ means brain signal of $S_{2}$ hearing $T_{11}$.
The line connecting two symbols indicates they are related to one sample in dataset. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it indicates that subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our data splitting method. As the splitting of validation set is the same as test set, we only consider training set and test set in this section for simplicity.
We will use method (a), (b), (c), (d), (e) to represent five current dataset splitting methods in the rest of the paper. Method (a) splits the dataset according to subjects, which can be described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|S_{i}\notin S_{val},S_{test}\}$ (5)
for training set. Method (b) splits the dataset according to tasks, which is described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|M_{k}\notin S_{val},S_{test}\}$ (6)
for training set. Method (c), (d), and (e) all split the dataset according to brain signal frames
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}.$ (7)
However, there are slight differences between these three methods. Method (c) views all the brain signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then unions all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames, while method (e) picks consecutive signal frames.
Type Method Narratives / ZuCo Average
seed1111 seed2222 seed3333 seed4444
@@ 3.2 Current Splitting Methods
(b) 6.73 / - 6.32 / - 7.7 / - 17.93 / - 9.67 / -
(c) 12.55 / 12.52 12.52 / 12.55 12.48 / 12.48 12.44 /12.46 12.50 / 12.50
(d) 12.81 / 12.60 12.8 / 12.58 12.78 / 12.56 12.79 / 12.61 12.795 / 12.59
(e) 12.28 / - 12.27 / - 12.26 / - 12.27 / - 12.27 / -
(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
[…]
(b) 0.00 / - 0.00 / - 0.00 / - 0.00 / - 0.00 / -
(c) 100.00 / 13.21 100.00 / 13.06 100.00 / 12.91 100.00 / 13.1 100.00 / 13.07
(d) 99.93 / 0.00 99.81 / 0.00 99.54 / 0.00 99.99 / 0.00 99.82 / 0.00
(e) 9.19 / - 9.31 / - 9.36 / - 9.29 / - 9.29 / -
(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
Table 1: Results of Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR).
We first point out the data leakage problems in current splitting methods through the analysis of training set and test set composition. Specifically, following the definition in subsection 3.1, two kinds of data leakage, brain signal leakage and text stimuli leakage, are defined. The data leakage situation of different methods can be reflected through visualization in Figure 2. Lines between $S_{i}$ and $M_{k}$ indicate brain signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation. If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set. Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two kinds of brain signal too. Since fMRI signals need to be sampled continuously with a certain length $L$, one sample shown in Figure 2 (e.g. $S_{1}$-$M_{1}$-$T_{11}$-$F_{11}$) is actually the first part of one fMRI sample, with $L-1$ continuous brain signal frames following (e.g. $F_{12},F_{13},\ldots$). In this sense, brain signal leakage doesn’t exist in method (a) for EEG, but method (a) suffers from text stimuli leakage. Text stimuli leakage does not exist in method (b) but brain signal leakage exists. Method (c) suffers from both brain signal leakage and text stimuli leakage. Method (d) and (e) are the same to EEG, with brain signal leakage. For fMRI, the situation of data leakage for different methods is similar to EEG, except for method (d) and (e), which are the same for EEG but actually different for fMRI. Method (d) suffers from both brain signal leakage and text stimuli leakage while in method (e) text stimuli leakage happens in the overlap between training samples and test samples.
3.3 Cross-Subject Splitting Criterion
@@
To eliminate data leakage from both brain signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2. Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two data splitting methods. As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pair in the dataset. $N$ is the total number of subjects and $M$ is the total number of unique text stimuli. We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject. As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 4, some edges satisfying zero data leakage condition are not included in the graph. We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
The sample in fMRI dataset is formatted in $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$. If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set. We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data leakage. Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm. More details and pseudo-code are available in Appendix B.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
@@
4 Experimental Settings
We test some SOTA brain-to-text decoding models on two popular cognitive datasets Narratives (Nastase et al., 2021) and ZuCo (Hollenstein et al., 2018). Dataset details are introduced in Appendix A. Comprehensive experiments are conducted to evidence the existence of the following phenomena: (1) Brain signals and text stimuli in test set leak into training set in all current dataset splitting methods. (2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data leakage. Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d) for EEG.
4.1 Implementation
We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except all the datasets are split to the ratio of 8:1:1 for fair comparison. All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs. More details are shown in Appendix A.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
@@ 4.1 Implementation
4.2 Data Leakage Metrics
We have analyzed two kinds of data leakage: brain signal leakage and text stimuli leakage in section 3. In this part, we will quantify two kinds of data leakage through experiments.
To better illustrate the extent of data leakage of different data splitting methods, we design two novel evaluation metrics named Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR) for detecting brain signal leakage and text stimuli leakage. Note that the situation for validation set is the same as test set, so we only consider test set in experiments. BSLR indicates the average percentage of each subject’s brain signals in test set appearing in training set, which could be formulated as
$\frac{1}{N}\sum_{i=1}^{N}\min(1,\frac{|\{F_{ij}|F_{ij}\in S_{test}\cap S_{train}\}|}{|\{F_{ij}|F_{ij}\in S_{train}\}|})$ (8)
where $N$ stands for the total number of subjects in test set. ||$\cdot$|| ⋅ | stands for the cardinality of a set. Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than one.
The definition of TSLR is somewhat different for EEG signal and fMRI signal. As to EEG signal where brain signals are sampled continuously, it’s easy to match certain text stimuli with corresponding signals. Its TSLR is similar to BSLR, which indicates the average percentage of certain text in test set appearing in training set. TSLR for EEG data can be calculated through
$\frac{1}{M}\sum_{j=1}^{M}\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{train}\}|})$ (9)
where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject.
The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to text. Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding text segments as training samples. As a result, we consider the average percentage of the same text segments in test set appearing in training set as TSLR for fMRI signal. It can be formulated as
$\frac{1}{M}\sum_{j=1}^{M}\tau\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|S_{test}|\times L}$ (10)
where $\tau=0$ if $\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}=\emptyset$ else
$\tau=\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}).$ (11)
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
Narratives UniCoRN 22.83 5.69 1.43 0.48 15.55 24.80 19.04
ZuCo UniCoRN 23.32 7.78 3.01 1.09 18.47 20.00 17.92
EEG2Text 24.49 7.49 2.28 0.62 23.98 23.95 25.74
Table 4: A fair benchmark for evaluating brain-to-text decoding.
5 Results and Analysis
5.1 Verification for Data Leakage
We test current data splitting methods and our method on fMRI dataset “Narratives” and EEG dataset ZuCo. Considering the influence of randomness in splitting, we randomly select four seeds for experiments. The results are shown in Table 1, and are consistent with theoretical analysis. For fMRI, current methods apart from method (a) suffer from brain signal leakage, while method (a) has serious text stimuli leakage. Method (b) gets no text stimuli leakage but has slight brain signal leakage. The situation for EEG is similar to that of fMRI. Apart from our proposed method (f), there is no way to achieve zero brain signal leakage and text stimuli leakage at the same time.
5.2 Damage of Data Leakage
Brain signal leakage and text stimuli leakage will damage brain-to-text decoding models from both the encoder side and decoder side.
Effect on Encoder
The encoder of current brain-to-text decoding models can be trained in two ways: either jointly with the decoder or solely pre-trained through a reconstruction task. In the former end-to-end training scenario, it is hard to evaluate encoder performance separately. So we mainly focus on the latter, in which case the encoder is pre-trained through an encoder-decoder framework to reconstruct input brain signals. The decoder here does not refer to the decoder for text generation. It is similar to the structure of the encoder and will be abandoned once the encoder is pre-trained. Since a proper evaluation index of the encoder’s representation ability is missing, validation loss is applied to measure the effect of data leakage.
We test different splitting methods on two cognitive datasets. The validation loss of encoder is shown in Figure 4. For fMRI, influenced by leakage of brain signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs. The encoder is actually overfitting and degrading. For method (a) and (f) without brain signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning. For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of brain signal leakage or not. We think the poor spatial resolution of EEG signal might lead to this phenomenon.
Effect on Decoder
All SOTA models choose the pre-trained language model BART (Lewis et al., 2020) as decoder. The powerful model is able to achieve fluent open vocabulary text generation. However, if data leakage occurs, due to the auto-regressive generation that calculates the probability of current token based on all previous tokens, the decoder will generate seen text given the first few words, and fail to generalize to unseen text.

### v4

@@ Appendix A Related Work
Appendix B Implementation Details
@@
We perform the same filtering steps to “Narratives” dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022). In BSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively. In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively. The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper. In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs. Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size. Sample length $L$ is set as 10 for all experiments related to fMRI. For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs. UniCoRN model is trained with the same settings as fMRI data.
Appendix C Cross-Subject Data Splitting in Practice
@@
We present the pseudo-code of two dataset splitting methods for EEG and fMRI signal. We only consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ instead of a $4$-partite graph in real practice. For EEG signal, $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$. While for fMRI signal, $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{K}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$. $N,M,K$ indicate the total number of subjects, text segments and stories. We assert $M>N$ for EEG dataset and $K<N$ for fMRI dataset, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text segment or story is listened by at least one subject. As shown in step 1 of Figure 4, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then following step 2, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 3, we extend each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ by adding edges without data leakage.
The main difference of splitting methods for EEG and fMRI lies in how $\mathcal{G}_{2}$ is generated. We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to generate $\mathcal{G}_{2}$. Specifically, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$. In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$. All assertions are based on real cognitive datasets. One more thing to notice is that in Line 14 of both pseudo-code, the loop indicates extending training set, validation set, and test set respectively. So the names of variable should be alternated in the repeat loop and the displayed part in pseudo-code is a case example of extending training set. We write it in this way for simplicity of expression.
Appendix D Supplementary Proof
Definition D.1.
An directed multigraph $\mathcal{G}$ is a type of graph which is permitted to have multiple edges between two vertices. When the edges own identity, $\mathcal{G}$ can be written as $\mathcal{G}=(\mathcal{V},\mathcal{E},f)$, where $f:\mathcal{E}\rightarrow\mathcal{V}\times\mathcal{V}$ is an incidence function that maps each edge to a pair of vertices.
Definition D.2.
A $k$-partite graph $\mathcal{G}$ is a type of graph that can be divided into $k$ distinct independent sets such that no two vertices in the same set are connected. $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{V}_{1}\cup\mathcal{V}_{2}\cup\cdots\cup\mathcal{V}_{k}$ and $\forall i\neq j,\mathcal{V}_{i}\cap\mathcal{V}_{j}=\emptyset$.
Notation D.3.
$\otimes$ is a Cartesian product-like operator. $X\otimes Y=\{(x,y)|x\in X,y\in Y,$ there exists relationship between $x$ and $y$ in dataset}\}. It’s designed to describe the connectivity among $\mathcal{S},\mathcal{M},\mathcal{T},\mathcal{F}$. For example, edges in $\mathcal{S}\otimes\mathcal{M}$ indicates certain subjects are stimulated by certain stories as described in dataset.
Definition D.4.
The training set for cross-subject brain-to-text decoding should be formatted in $\mathcal{G}_{train}=\mathcal{S}_{train}\otimes\mathcal{M}\otimes\mathcal{T}_{train}\otimes\mathcal{F}_{train}$, where $\mathcal{S}_{train}=\{S_{i}|\forall S_{i}^{\prime}\in\mathcal{S}_{test},S_{i}\neq S_{i}^{\prime}\}$; $\mathcal{F}_{train}=\{F_{ijk}|i\in I\}$, $I=\{i|\forall j,\forall k,F_{ijk}\notin\mathcal{F}_{test}\}$; $\mathcal{T}_{train}=\{T_{kj}|\forall T_{kj}^{\prime}\in\mathcal{T}_{test},T_{kj}\neq T_{kj}^{\prime}\}$.
Why a method without brain signal leakage and text stimuli leakage must satisfy cross-subject brain-to-text decoding criterion
Training set $\mathcal{G}_{train}$ without brain signal leakage and text stimuli leakage is formatted in
$\mathcal{G}_{train}$ $=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (21)
$\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},$
$S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}$
$=\mathcal{S}_{train}\otimes\mathcal{M}\otimes\mathcal{T}_{train}\otimes\mathcal{F}$
where $\mathcal{S}_{train}=\{S_{i}|\forall S_{i}^{\prime}\in\mathcal{S}_{test},S_{i}\neq S_{i}^{\prime}\},\mathcal{T}_{train}=\{T_{kj}|\forall T_{kj}^{\prime}\in\mathcal{T}_{test},T_{kj}\neq T_{kj}^{\prime}\}$. Since $F_{ijk}\in\mathcal{F}$ indicates brain signal of subject $S_{i}$ stimulated by text segment $T_{kj}$, and given the definition of operator $\otimes$, $\mathcal{F}$ is determined when $\mathcal{S}$ and $\mathcal{T}$ are specified, which is
$\mathcal{F}$ $=\{F_{ijk}|i\in I,kj\in J\},$ (22)
$I$ $=\{i|S_{i}\in\mathcal{S}_{train}\},$
$J$ $=\{kj|T_{kj}\in\mathcal{T}_{train}\}.$
$\mathcal{F}$ can also be written as $\mathcal{F}=\{F_{ijk}|i\in I\}$, $I=\{i|\forall j,\forall k,F_{ijk}\notin\mathcal{F}_{test}\}$, which is equal to Definition D.4.
Why the proposed splitting method satisfy zero data leakage
Take the splitting method for EEG signal as example, the training set and test set after step 1 and step 2 already satisfy
$\mathcal{G}_{train}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (23)
$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}$
$\mathcal{G}_{test}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (24)
$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{train},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}$
So we only need to prove expanded graph $\mathcal{G}_{train\_exp}^{\prime}$ and $\mathcal{G}_{test\_exp}^{\prime}$ satisfy zero data leakage, which is obvious from Equation 13 and 14.
Why we must discard samples to ensure no data leakage
If $\mathcal{G}_{train}\cup\mathcal{G}_{test}=\mathcal{G}_{\mathcal{D}}$, suppose $\forall(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{train}$, $(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}$. For $f(\mathcal{E})=(M_{k},T_{kj})$, $f(\mathcal{E}^{\prime})=(M_{k}^{\prime},T_{kj}^{\prime})$, $T_{kj}\neq T_{kj}^{\prime}$:
  • If $M_{k}=M_{k}^{\prime}$, then there must exist a subject $S_{i}=S_{i}^{\prime}$ such that he is stimulated by the whole stories.
  • If $M_{k}\neq M_{k}^{\prime}$, then there must exist a subject $S_{i}=S_{i}^{\prime}$ such that he is stimulated by two different stories.
As a result, if $\mathcal{G}_{train}\cup\mathcal{G}_{test}=\mathcal{G}_{\mathcal{D}}$, then $\exists(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{train}$, $(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test}$, $s.t.$ $S_{i}=S_{i}^{\prime}$ or $T_{kj}=T_{kj}^{\prime}$. Some samples must be discarded to ensure no data leakage.
Dataset Model Method With Teacher-Forcing / Without Teacher-Forcing
BLEU-1 BLEU-2 BLEU-3 ROUGE1-F
Narratives UniCoRN (a) 49.56 / 33.82 30.49 / 9.44 21.07 / 4.89 40.65 / 21.87
(b) 26.37 / 19.68 7.50 / 5.50 2.48 / 1.76 19.62 / 16.86
(c) 50.24 / 29.87 30.83 / 8.35 21.23 / 3.21 41.01 / 19.03
(d) 49.63 / 30.20 30.29 / 8.78 20.85 / 3.65 41.03 / 19.42
(e) 28.94 / 21.46 9.39 / 6.24 4.07 / 1.83 19.49 / 17.39
(f) 22.83 / 16.85 5.69 / 4.24 1.43 / 0.65 19.04 / 15.34
ZuCo UniCoRN (a) 58.09 / 19.47 49.23 / 7.69 43.23 / 2.97 67.50 / 17.25
(c) 52.30 / 19.70 42.89 / 7.54 36.80 / 2.93 67.29 / 17.37
(d) 50.02 / 22.02 43.53 / 8.28 32.71 / 3.15 67.33 / 18.33
(f) 23.32 / 14.02 7.78 / 2.57 3.01 / 0.82 17.92 / 11.95
EEG2Text (a) 51.22 / 21.99 33.83 / 7.42 22.99 / 2.94 46.58 / 17.78
(c) 53.83 / 20.41 38.99 / 7.25 29.57 / 2.48 53.56 / 17.32
(d) 53.92 / 19.80 41.06 / 7.46 23.12 / 3.06 49.38 / 17.29
(f) 24.49 / 13.52 7.49 / 2.86 2.28 / 0.78 25.74 / 11.20
Table 6: Performance of brain-to-text decoding models under different splitting methods with teacher-forcing and without teacher-forcing. The green mark and red mark denotes a method without and with text stimuli leakage correspondingly.
Model Epoch+lr+Method BLEU-N ($\%$) ROUGE-1 ($\%$)
$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
@@ Appendix D Supplementary Proof
Model Epoch+lr+Method BLEU-N ($\%$) ROUGE-1 ($\%$)
$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
@@
Appendix E Supplementary Experiment
The influence of teacher-forcing
Jo et al. (2024) pointed out that previous methods applied teacher-forcing during generation, which has the tendency to cause overfitting problems. Therefore, we conduct experiments on models without teacher-forcing to ensure that the affection of data leakage is not influenced by teacher-forcing. Results are shown in Table 6. Without the influence of teacher-forcing decoding, splitting methods with data leakage will still lead to overestimation of model performance.
Experiments on longer training epochs with smaller learning rate